# On annihilator multiplication modules

### v1

@@ 3. Characterizations of some special rings/modules
$(i)\Rightarrow(ii):\$Suppose that AA\ is a reduced ring and every faithful $A$-module is an annihilator multiplication module. Let $I$ be an ideal of $A$, and put $E=A\oplus\left(A/I\right).$ Note that EE\ is a faithful multiplication module, by assumption, EE\ is an annihilator multiplication module. This implies that $ann((0,\overline{1}))=I=ann(JE)=ann(J)$ for some finitely generated ideal $J$ of A.A.\ Then there exist $x_{1},x_{2},\ldots,x_{n}\in J$ such that $J={\textstyle\sum\limits_{i=1}^{n}}Ax_{i}.\$This implies that $I={\textstyle\bigcap\limits_{i=1}^{n}}ann(x_{i}).\$Now, take a prime ideal $P$ of A.A.\ Then we have $P={\textstyle\bigcap\limits_{i=1}^{n}}ann(y_{i})$ for some $y_{1},y_{2},\ldots,y_{n}\in A.\$Since PP\ is prime, we conclude that $P=ann(y_{i})$ for some $y_{i}\in A.\$As AA\ is a reduced ring, by [16, Lemma 2.8], PP\ is a minimal prime ideal of A.A.\ Thus, every prime ideal of AA\ is maximal, so AA\ is a von Neumann regular ring. Now, we will show that AA\ is a principal ideal ring. Let II\ be an ideal of A.A.\ Then by above, $I={\textstyle\bigcap\limits_{i=1}^{n}}ann(x_{i}).\$Since AA\ is a von Neumann regular ring, every principal ideal $(x_{i})=(b_{i})\$is generated by an idempotent $b_{i}\in R.\$Thus, we conclude that $ann(x_{i})=(1-b_{i}),\$where $1-b_{i}\$is again an idempotent of A.A.\ Since $(b)\cap(c)=(bc)\$for every two idempotents $b,c\in A,\$we obtain
@@
$\left(ii\right)\Rightarrow\left(iii\right):\$Let AA\ be a principal ideal ring and von Neumann regular ring. Then every prime ideal is maximal and AA\ is reduced. Since AA\ is principal ideal ring, $\sqrt{0}=(0)=Jac(A).$ Since AA\ is Noetherian and zero dimensional, it is Artinian, and so AA\ has only finitely many maximal ideals $Q_{1},Q_{2},\ldots,Q_{n}.\$This gives $Q_{1}\cap Q_{2}\cap\cdots\cap Q_{n}=(0).\$Then by Chinese Remainder Theorem, we have $A=A/(0)\cong\left(A/Q_{1}\right)\oplus\left(A/Q_{2}\right)\oplus\cdots\oplus\left(A/Q_{n}\right)$, where $A/Q_{i}\$is a field.
$\left(iii\right)\Rightarrow(ii):$ It is clear.
@@
Let $E$ be an $A$-module. An ideal $P$ of $A$ is defined as an associated prime of $E$ if $P$ is a prime ideal and $P=ann(e)$ for some $0\neq e\in E$. The set of all associated primes of $E$ is denoted by $Ass_{A}(E)$ [25]. In this context, $Ass(A)$ denotes $Ass_{A}(A)$. Associated prime ideals correspond to the irreducible components underlying a module or ideal. These ideals play a fundamental role in primary decomposition, establish connections between algebraic structures and the geometry of varieties, and are crucial in the study of dimension and depth theory. The following section presents the relationships between $Ass_{A}(E)$ and $Ass(A)$ in terms of annihilator multiplication modules.
Let $E$ be an $A$-module. The set of all torsion elements of $E$ is denoted by $T(E)=\{e\in M:ann(e)\neq 0\}$. According to [6], $E$ is classified as a torsion module if $T(E)=E$; otherwise, $E$ is referred to as a non-torsion module.
@@
(i) Let EE\ be an annihilator multiplication module faithful module. Then $Ass_{A}(E)\subseteq Ass(A).$
(ii) Let EE\ be a non torsion module and annihilator multiplication module. Then $Ass(A)=Ass_{A}(E).\$
@@
(i):(i):\ Let EE\ be an annihilator multiplication module faithful module. Choose $P\in Ass_{A}(E).\$Then PP\ is a prime ideal and there exists $0\neq e\in E$ such that P=ann(e).P=ann(e).\ Since EE\ is an annihilator multiplication module, there exists a finitely generated ideal II\ of AA\ such that ann(e)=ann(IE).ann(e)=ann(IE).\ Then we can write $I={\textstyle\sum\limits_{i=1}^{n}}Aa_{i}$ for some $a_{1},a_{2},\ldots,a_{n}\in I.\$This implies that
@@
(ii):(ii):\ Since every non torsion module is faithful, by (i),(i),\ we have $Ass_{A}(E)\subseteq Ass(A).$ For the reverse inclusion, choose $P\in Ass(A).\$Then there exists $0\neq a\in A$ such that P=ann(a)P=ann(a)\ is a prime ideal of A.A.\ Since EE\ is a non torsion module, there exists $0\neq e\in E\$such that $ann(e)=(0)$ which implies that $P=ann(a)=ann(ae),$ that is, $P\in Ass_{A}(E).\$Then we have the equiality $Ass(A)=Ass_{A}(E).\$
Conflict of Interest. The corresponding author states that there is no conflict of interest.

### v2

@@ 3. Characterizations of some special rings/modules
$(i)\Rightarrow(ii):\$Suppose that AA\ is a reduced ring and every faithful $A$-module is an annihilator multiplication module. Let $I$ be an ideal of $A$, and put $E=A\oplus\left(A/I\right).$ Note that EE\ is a faithful module, by the assumption, EE\ is an annihilator multiplication module. This implies that $ann((0,\overline{1}))=I=ann(JE)=ann(J)$ for some finitely generated ideal $J$ of A.A.\ Then there exist $x_{1},x_{2},\ldots,x_{n}\in J$ such that $J={\textstyle\sum\limits_{i=1}^{n}}Ax_{i}.\$This implies that $I={\textstyle\bigcap\limits_{i=1}^{n}}ann(x_{i}).\$Now, take a prime ideal $P$ of A.A.\ Then we have $P={\textstyle\bigcap\limits_{i=1}^{n}}ann(y_{i})$ for some $y_{1},y_{2},\ldots,y_{n}\in A.\$Since PP\ is prime, we conclude that $P=ann(y_{i})$ for some $y_{i}\in A.\$As AA\ is a reduced ring, by [22, Lemma 2.8], PP\ is a minimal prime ideal of A.A.\ Thus, every prime ideal of AA\ is maximal, so AA\ is a von Neumann regular ring. Now, we will show that AA\ is a principal ideal ring. Let II\ be an ideal of A.A.\ Then by above, $I={\textstyle\bigcap\limits_{i=1}^{n}}ann(x_{i}).\$Since AA\ is a von Neumann regular ring, every principal ideal $(x_{i})=(b_{i})\$is generated by an idempotent $b_{i}\in A.\$Thus, we conclude that $ann(x_{i})=(1-b_{i}),\$where $1-b_{i}\$is again an idempotent of A.A.\ Since $(b)\cap(c)=(bc)\$for every two idempotents $b,c\in A,\$we obtain
@@
$\left(ii\right)\Rightarrow\left(iii\right):$ It is straightforward.
$\left(iii\right)\Leftrightarrow\left(iv\right):$ Follows from [30, Corollary 2.6].
$\left(iv\right)\Rightarrow\left(ii\right):$ It is clear.
@@
Let $E$ be an $A$-module. An ideal $P$ of $A$ is defined as an associated prime of $E$ if $P$ is a prime ideal and $P=ann(e)$ for some $0\neq e\in E$. The set of all associated primes of $E$ is denoted by $Ass_{A}(E)$ [33]. In this context, $Ass(A)$ denotes $Ass_{A}(A)$. Associated prime ideals correspond to the irreducible components underlying a module or ideal. These ideals play a fundamental role in primary decomposition, establish connections between algebraic structures and the geometry of varieties, and are crucial in the study of dimension and depth theory. The following result presents the relationships between $Ass_{A}(E)$ and $Ass(A)$ in terms of annihilator multiplication modules.
Let $E$ be an $A$-module. The set of all torsion elements of $E$ is denoted by $T(E)=\{e\in E:ann(e)\neq 0\}$. According to [4], $E$ is said to be a torsion module if $T(E)=E$; otherwise, $E$ is called a non-torsion module.
@@
(i) Let EE\ be an annihilator multiplication faithful $A$-module. Then $Ass_{A}(E)\subseteq Ass(A).$
(ii) Let EE\ be an annihilator multiplication $A$-module which is also non-torsion. Then $Ass(A)=Ass_{A}(E).\$
@@
(i):(i):\ Let EE\ be an annihilator multiplication faithful module. Choose $P\in Ass_{A}(E).\$Then PP\ is a prime ideal and there exists $0\neq e\in E$ such that P=ann(e).P=ann(e).\ Since EE\ is an annihilator multiplication module, there exists a finitely generated ideal II\ of AA\ such that ann(e)=ann(IE).ann(e)=ann(IE).\ Then we can write $I={\textstyle\sum\limits_{i=1}^{n}}Aa_{i}$ for some $a_{1},a_{2},\ldots,a_{n}\in I.\$This implies that
@@
(ii):(ii):\ Since every non-torsion module is faithful, by (i),(i),\ we have $Ass_{A}(E)\subseteq Ass(A).$ For the reverse inclusion, choose $P\in Ass(A).\$Then there exists $0\neq a\in A$ such that P=ann(a)P=ann(a)\ is a prime ideal of A.A.\ Since EE\ is a non torsion module, there exists $0\neq e\in E\$such that $ann(e)=(0)$ which implies that $P=ann(a)=ann(ae),$ that is, $P\in Ass_{A}(E).\$Then we have the equiality $Ass(A)=Ass_{A}(E).\$
The following example demonstrates why the conditions of ”annihilator multiplication” or ”faithful” are necessary for Theorem 6 (i) to hold. Also, it shows that the conditions ”annihilator multiplication module” or ”non-torsion module” in Theorem 6 (ii) can not be removed.
Example 8.
  • (i)
    Consider the $\mathbb{Z}$-module $E=\mathbb{Z}_{p}\oplus\mathbb{Z}$, where $p$ is a prime number. Then $E$ is not an annihilator multiplication module by Example 3. Also it is easy to check that $E$ is a faithful (even, a non-torsion) $\mathbb{Z}$-module and $Ass_{\mathbb{Z}}(E)=\{(0),p\mathbb{Z}\}\nsubseteq Ass(\mathbb{Z})=\{(0)\}$.
  • (ii)
    Consider the $\mathbb{Z}$-module $E=\mathbb{Z}_{p}$, where $p$ is a prime number. Then $E$ is a multiplication module, so by Example 1, it is an annihilator multiplication module. Also, $E$ is not a faithful module since $ann(E)=p\mathbb{Z}\neq 0$. However, $Ass_{\mathbb{Z}}(E)=\{p\mathbb{Z}\}$ and $Ass(\mathbb{Z})=\{(0)\}$ are not comparable.
  • (iii)
    Consider the $\mathbb{Z}$-module $E=\bigoplus_{i=1}^{\infty}\mathbb{Z}_{p}$ for some prime number pp. Then by Proposition 5 and Example 1, $E$ is an annihilator multiplication module. However, $E$ is a torsion module. Also, one can easily verify that $Ass(\mathbb{Z})=\{(0)\}$ and $Ass_{\mathbb{Z}}(E)=\{p\mathbb{Z}\}$ are distinct.
While the examples provided in Example 8 highlight the significance of the annihilator multiplication property in the study of associated prime ideals, the necessity of the ”non-torsion” condition in Theorem 6 (ii) remains an intriguing question. Specifically, if $E$ is assumed to be a faithful torsion annihilator multiplication module, it remains unknown whether the equality $Ass_{A}(E)=Ass(A)$ holds without any additional assumptions, as we have yet to find a counterexample. Consequently, we leave the validity of this assertion as an open problem for the reader.
Question 1.
Let $E$ be an annihilator multiplication faithful torsion $A$-module. Is it true that $Ass(A)=Ass_{A}(E)$ ?
Declarations
Conflict of Interest: The authors declare that they have no conflict of interest.
Data Availability Statement: Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.